OpenStudy (lena772):

At 1000 K, Kc = 2.11E-2. If one starts with 1.00 mole graphite and 1.00 mole CO2 (g) in a 40.0 L vessel, what is the equilibrium concentration of CO (g)?

OpenStudy (lena772):

I caclulated ad got x=0.021644, but when I plugged that in to find the conc. of CO it was wrong.

OpenStudy (lena772):

Cs, graphite + CO2 (g) ⇌ 2 CO (g)

OpenStudy (lena772):

@Abhisar

OpenStudy (jfraser):

if you included the "concentration" of the solid graphite, you shouldn't

OpenStudy (lena772):

i did. do i sassume 2 moles of CO as well because it doesn't explicitly say that but its 1to1to2

OpenStudy (jfraser):

the balanced reaction will create 2 moles of CO, you're correct. But you don't have 2 moles of CO at equilibrium

OpenStudy (jfraser):

remember that in equilibrium, only \(some\) of the reactants combine to form \(some\) of the products Use an ICE table approach and set the unknown concentrations equal to the Kc value

OpenStudy (lena772):

I mean is the CO value for moles at initial 2? b/c then you would say 2mol/40.0L of soln to get the initial conc. of CO? Which is 0.05

OpenStudy (jfraser):

your initial value of CO in the flask is zero. You need to find the concentrations \(at\) \(equiliubrium\) \[K_C = \frac{[CO]^2}{[CO_2]}\]

OpenStudy (lena772):

but kc is given

OpenStudy (jfraser):

KC is given, but the concentrations are missing. that's what you're solving for

OpenStudy (jfraser):

the starting concentration of \(CO_2\) is 0.025M, correct?

OpenStudy (lena772):

yes

OpenStudy (lena772):

but 0^2/0.025 is 0...

OpenStudy (jfraser):

the \(equilibrium\) concentration of \(CO_2\) will be (0.025M - x)

OpenStudy (lena772):

yes

OpenStudy (jfraser):

the equilibrium concentration of \(CO\) is not zero

OpenStudy (jfraser):

as the \(CO_2\) is used, some CO is \(formed\)

OpenStudy (jfraser):

for every "x" amount of \(CO_2\) that is used, the amount of \(CO\) that is formed is \(twice\) that amount, because of the balanced reaction

OpenStudy (jfraser):

so the equilibrium concentration of \(CO\) is "2x"

OpenStudy (lena772):

wouldn't it be -2x cause Q<Kc

OpenStudy (jfraser):

\(CO\) is the product \(formed\), so it's amount is "2x"

OpenStudy (jfraser):

the equilibrium equation then looks like this

OpenStudy (lena772):

would it be 2.11e-2=(2x)^2/(0.025+x)?

OpenStudy (jfraser):

\[K_C = \frac{[CO]^2}{[CO_2]} =\frac{[2x]^2}{[0.025-x]} = 0.0211\]

OpenStudy (jfraser):

the \(CO_2\) is \(used \space up\) so its equilibrium concentration must be less than what you started with

OpenStudy (lena772):

I just thought when Q<Kc the equilibrium shifted left :/

OpenStudy (jfraser):

it does, but that's not what this question is asking you for

OpenStudy (jfraser):

it tells you that you start with graphite \(C(s)\) and \(CO_2\). there is no \(CO\) formed yet, so the reaction \(must\) proceed forward

OpenStudy (lena772):

0.01829034 M = CO at equilibrium?

OpenStudy (lena772):

No i didn't. I get x=-0.0144202, x=0.00914517. I multiplied the second x value by 2, and that's what I got...

OpenStudy (lena772):

Multiplying the first one gives a negative, and we can't have negative concentration

OpenStudy (jfraser):

did you solve it quadratically, or take the lazy shortcut?

OpenStudy (lena772):

i graphed it.

OpenStudy (jfraser):

what did you plug in for the coefficients?

OpenStudy (lena772):

2.11e-2=((2x)^2)/(0.025-x)

OpenStudy (jfraser):

I don't normally solve them that way, but it's probably close enough. I got 0.0229M

OpenStudy (lena772):

are you good at lab reports?

OpenStudy (lena772):

@JFraser

OpenStudy (jfraser):

usually, but what things you should put in it really depends on what your teacher has asked for